# Show Less, Instruct More: Enriching Prompts with Definitions and Guidelines for Zero-Shot NER

### v2

@@ Abstract
Recently, several specialized instruction-tuned Large Language Models (LLMs) for Named Entity Recognition (NER) have emerged. Compared to traditional NER approaches, these models have strong generalization capabilities. Existing LLMs mainly focus on zero-shot NER in out-of-domain distributions, being fine-tuned on an extensive number of entity classes that often highly or completely overlap with test sets. In this work instead, we propose SLIMER, an approach designed to tackle never-seen-before named entity tags by instructing the model on fewer examples, and by leveraging a prompt enriched with definition and guidelines. Experiments demonstrate that definition and guidelines yield better performance, faster and more robust learning, particularly when labelling unseen Named Entities. Furthermore, SLIMER performs comparably to state-of-the-art approaches in out-of-domain zero-shot NER, while being trained on a reduced tag set.
@@
Figure 1: SLIMER’s instruction-tuning prompt. Dedicated entity definition and guidelines steer the model generation.
@@ 1 Introduction
Named Entity Recognition (NER) is a crucial problem in Natural Language Processing (NLP), usually being a key component in Information Extraction pipelines. Traditional methods frame NER into a sequence labeling task Li et al. (2020), where models are specialized on a narrow domain and a pre-defined label set, thus lacking generalization capabilities outside the downstream task at hand.
On the contrary, Large Language Models (LLMs) have demonstrated strong zero-shot capabilities. Through the use of cleverly designed prompts, models like GPT-3 can tackle NER via In-Context Learning Radford et al. (2019); Brown et al. (2020). However, smaller (encoder-only) LMs trained on specific NER tasks may still outperform LLMs (Wang et al., 2023a; Ye et al., 2023; Zhou et al., 2023). To this end, Instruction-Tuning of LLMs has emerged as an effective method to improve their performance Wei et al. (2022); Chung et al. (2022); Wang et al. (2022b). There exist several works in the literature based on instruction-tuning for NER, such as InstructUIE Wang et al. (2023b), UniNER Zhou et al. (2023), GoLLIE Sainz et al. (2024) and GNER Ding et al. (2024).
@@
Such dedicated LLMs have the ability to perform zero-shot NER on heterogeneous input domains and a multitude of possibly never-seen-before NEs. Existing works mainly focus on zero-shot NER in Out-Of-Distribution (OOD) input domains, while fine-tuning on an extensive number of entity classes that often highly or completely overlap between the training and test sets. Consequently, the problem of tagging unseen named entities has been little investigated, with GoLLIE as the only exception.
In this work instead, we tackle both scenarios by carefully selecting a training set with limited degree of overlap with test data. To facilitate effective zero-shot capabilities on such novel NEs, similar to GoLLIE, we steer the model with annotation guidelines. Unlike GoLLIE, we drop their code-based representation in favour of a more natural instruction that includes a definition and some guidelines for the category being extracted. This simplifies the prompt design and it allows us to automatically generate synthetic definition and guidelines by means of another LLM.
We study how increasing both the number of training samples and the number of unique Named Entity (NE) categories affect the generalization capabilities of LLMs, with or without the support of definitions and guidelines. By using fewer training samples from a reduced number of distinct named entity tags, combined with prompts enriched of definition and guidelines, we name our approach SLIMER: Show Less, Instruct More - Entity Recognition [^1] https://huggingface.co/expertai/SLIMER.
Experiments were conducted on two standard NER benchmarks for OOD, MIT (Liu et al., 2013) and CrossNER (Liu et al., 2021). Additionally, we assessed performance on never-seen-before NEs on BUSTER Zugarini et al. (2023), a document-level NER dataset with entity tags novel to all the evaluated models.
Comparison of SLIMER with its baseline, i.e. the model devoid of definition and guidelines, reveals SLIMER’s deeper understanding, faster and more stable learning, and better zero-shot performance. Despite being trained on a fraction of the data, with little overlap between train and test named entity tags, SLIMER performs comparably against state-of-the-art instruction-tuning approaches, revealing stronger generalization capabilities when dealing with unseen named entities.
MIT CrossNER BUSTER TOT
@@
$0/12$ $1/8$ $4/13$ $4/11$ $4/12$ $4/8$ $4/16$ $0/6$ $21/86$
$0\%$ $13\%$ $31\%$ $36\%$ $33\%$ $50\%$ $25\%$ $0\%$ $24\%$
Table 2: Overlap (%) between the NEs seen by SLIMER in training and those present in MIT and CrossNER benchmarks. For UniNER, GLINER and GNER models this overlap is 100%, i.e. there are no unseen NEs in test. The NEs in BUSTER are novel to all models.
@@ 2 Related Work
Commonly employed machine-learning solutions frame NER into a “sequence labeling task”, where the goal is to assign a BIO label to each element in a given sequence Li et al. (2020). Fine-tuning BERT-family models (Devlin et al., 2018) for NER is a well established approach. While these models excel in supervised contexts, they have the severe limitation of being constrained to a predefined set of labels and inputs from limited domains, making it difficult to generalize across different contexts and on unseen named entities.
@@ 2.1 In-Context Learning
Radford et al. (2019) were among the first to explore the ability of LLMs to perform “multi-task learning”. In this setting, tasks are formulated as text-to-text problems and natural language instructions are prefixed in the input to prompt the model towards the task it has to solve. Building on this breakthrough, Brown et al. (2020) enhanced these zero-shot and few-shot multi-task capabilities, paving the way for what has been termed “In-Context Learning” or “Prompt Engineering”.
While LLMs have demonstrated impressive capabilities in zero-shot settings on various challenging tasks, endeavours to utilize LLMs for Information Extraction have been less promising Keraghel et al. (2024). Attempts to make use of LLMs (such as GPT) through clever prompt engineering have been conducted by Wang et al. (2023a) in their paper GPT-NER. In Ye et al. (2023), the authors compared several GPT models on various NLU tasks, including NER. The results highlighted a significant gap compared to supervised encoder-only approaches. Those results were further confirmed by Zhou et al. (2023) on other NER datasets.
@@ 2.2 Fine-Tuning for Zero-Shot NER
Parallel to prompt engineering LLMs, other approaches explored the way of instruction-tuning LLMs Wei et al. (2022); Chung et al. (2022); Wang et al. (2022b), often involving models of smaller size and focusing the training on a single task, for example NER. InstructUIE Wang et al. (2023b) is a encoder-decoder T5-11B model fine-tuned on supervised IE datasets (among which NER) phrased as text-to-text problems, where the instruction exhorts the model to return a list of tuples (text span, entity type), choosing between a provided list of categories. UniNER Zhou et al. (2023) consists in a decoder-only LLaMA model finetuned on a “conversation style template”. In inference, the model is prompted with the question “What describes NE in the text?” and a list of text spans that belong to the requested NE category is returned.
Based on this instruction-tuning approach, several other works have emerged, each based on a different instruction-tuning template, with the aim of further improving the performance of LLMs in both supervised and zero-shot NER. Table 1 provides an overview of some of the most significant current approaches, highlighting the adopted backbone architecture, the designed instruction-tuning template, and other key comparative features.
GoLLIE Sainz et al. (2024) is the first (and currently the only) that includes annotation guidelines in its prompt. In particular, the authors adopted a code-based representation encoding the NE labels as Python classes and providing the guidelines as doc-strings. So far, GoLLIE is the only approach that was also evaluated on unseen named entities tagging (see Table 1).
GNER Ding et al. (2024) rethinks the importance of negative instances (i.e., “O” tags in BIO labeling) and replaces the established entity-centric schema with a BIO-like generation, replicating the same input text along with token-by-token BIO labels. Despite the limitations on the input length and output parsing difficulties, their approach displays stronger boundaries detection and reduced classification indecision compared to the other approaches.
Finally, GLiNER Zaratiana et al. (2023) consists of a much smaller model based on an encoder-only and non-instruction-tuned architecture, which achieves a remarkable performance in both supervised and zero-shot NER.
@@ 3 SLIMER
This section presents our approach, SLIMER. First, we provide motivations for reducing the number of training samples. Then, we describe the instruction-tuning prompt. Finally, we discuss how to generate definitions and guidelines automatically by means of another LLM, such as ChatGPT.
@@ 3.3 Definition and Guidelines generation
To overcome this limitation, we exploited the OpenAI’s Chat-GPT APIs to automatically generate definition and guideliens. In particular, we designed the one-shot prompt template reported in Figure 2, which was used to query gpt-3.5-turbo-1106. An exemplary one-round user-system conversation is used to illustrate the desired output to the model. The three examples are randomly sampled for each NE from the dataset at hand. Thanks to such a prompt, all the generated definitions and guidelines exhibit a similar structure, with short and clear defining sentences, and with guidelines highlighting edge cases where to be cautious.
@@ 4 Experiments
In the experiments, we investigate the impact of definition and guidelines on zero-shot NER. We compare SLIMER against state-of-the-art models on out-of-distribution input domains and unseen entity types. Furthermore, we study how increasing the number of training samples affects performance.
@@ 4.1 Datasets
The LLMs have all been fine-tuned on a subset of PileNER-type (Zhou et al., 2023), and evaluated on three different benchmarks: MIT (Liu et al., 2013) and CrossNER (Liu et al., 2021) to assess out-of-distribution (OOD) performance, BUSTER (Zugarini et al., 2023) to measure performance on never-seen-before Named Entities.
@@
PileNER-type (Zhou et al., 2023) is a synthetic dataset comprising a large set of approximately 50 thousands examples, encompassing over 13 thousands different named entity types. We kept only a subset of them, considering only those NEs with at least 100 instances. From the remaining 455 named entity classes, we manually revised and merged together classes of identical types spelled differently, e.g. organisation and organization, and discarded some “catch-all” gpt-hallucinated labels (e.g. unknown, other, miscellaneous, general, entity type), thus further reducing to 423 different labels. Finally, to limit the overlap between training and test entity types, we excluded nearly all the categories present in the test datasets, with the exception of the standard NER tags: person, location, organization, country. Overall, we kept 391 distinct NEs. Percentages of overlap between train and test are reported in Table 2. The dataset composition, grouped by topic, is illustrated in Figure 3.
@@
MIT (Liu et al., 2013) and CrossNER (Liu et al., 2021) datasets have become de-facto the standard benchmark for zero-shot NER. We use them to compare SLIMER against existing state-of-the-art models on out-of-distribution (OOD) domains.
@@
We extend the evaluation beyond the MIT and CrossNER benchmarks by including BUSTER (Zugarini et al., 2023). The dataset is a document-level NER benchmark in financial domain. Both domain and named entity tags differ from what observed by all the models during instruction-tuning. Its significant differences from standard NER datasets make BUSTER a perfect benchmark for evaluating zero-shot performance on never-seen-before NEs.
Figure 3: The 391 Named Entities in the PileNER-type subset, grouped by macro topics. “misc” (not shown) groups the 26 NEs that do not fit into the defined topics.
@@ 4.2 Settings
Backbone and Training setup.
SLIMER is based on LLaMA-2 7B chat Touvron et al. (2023). Investigating how different families or model sizes affect results is outside the objectives of our work. In all the experiments the models were fine-tuned with LoRA (Hu et al., 2021) $r=8$, $\alpha=16$ for 10 epochs with early stopping, 32 batch size and learning rate initialized to $3.0\times 10^{-4}$ with cosine scheduler and a warm-up of 60 steps. Context length was set to 768, longer inputs were chunked.
@@
Figure 4: Micro and macro F1 scores of SLIMER and its baseline without D&G on MIT, CrossNER and BUSTER altogether, as we respectively increase the number of unique NEs (left) and the number of samples per NE (right) seen in training.

### v3

@@ Abstract
Recently, several specialized instruction-tuned Large Language Models (LLMs) for Named Entity Recognition (NER) have emerged. Compared to traditional NER approaches, these models have demonstrated strong generalization capabilities. Existing LLMs primarily focus on addressing zero-shot NER on Out-of-Domain inputs, while fine-tuning on an extensive number of entity classes that often highly or completely overlap with test sets. In this work instead, we propose SLIMER, an approach designed to tackle never-seen-before entity tags by instructing the model on fewer examples, and by leveraging a prompt enriched with definition and guidelines. Experiments demonstrate that definition and guidelines yield better performance, faster and more robust learning, particularly when labelling unseen named entities. Furthermore, SLIMER performs comparably to state-of-the-art approaches in out-of-domain zero-shot NER, while being trained in a more fair, though certainly more challenging, setting.
@@
Figure 1: SLIMER’s prompt. Dedicated entity definition and guidelines steer the model generation.
@@ 1 Introduction
Named Entity Recognition (NER) is a crucial problem in Natural Language Processing (NLP), usually being a key component in Information Extraction pipelines. Traditional methods frame NER into a sequence labeling task Li et al. (2020), where models are specialized on a narrow domain and a pre-defined label set, thus lacking generalization capabilities outside the downstream task at hand. On the contrary, Large Language Models (LLMs) have demonstrated strong zero-shot capabilities. Through cleverly designed prompts, models like GPT-3 can tackle NER via In-Context Learning Radford et al. (2019); Brown et al. (2020). However, smaller (encoder-only) LMs, trained on the specific NER task, may still outperform LLMs (Wang et al., 2023a; Ye et al., 2023; Zhou et al., 2023). To this end, Instruction-Tuning of LLMs has emerged as an effective method to improve their performance Wei et al. (2022); Chung et al. (2022); Wang et al. (2022b). In the literature, several works have explored Instruction-Tuning for NER, including InstructUIE Wang et al. (2023b), UniNER Zhou et al. (2023), GoLLIE Sainz et al. (2024) and GNER Ding et al. (2024).
@@
Such dedicated LLMs have the ability to perform zero-shot NER on heterogeneous input domains and a multitude of possibly never-seen-before NEs. Existing works mainly focus on zero-shot NER on Out-Of-Domain (OOD) inputs, while fine-tuning on an extensive number of entity classes that often highly or completely overlap between the training and test sets (see Table 2). Consequently, the problem of tagging unseen named entities has been little investigated, with GoLLIE as the only exception.
In this work instead, we tackle both scenarios by carefully selecting a training set with minimal degree of class-overlap with test data. To facilitate effective zero-shot capabilities on such novel NEs, we guide the model with dedicated definition and guidelines for the category to be annotated. By using fewer training samples from a reduced number of distinct named entity tags, combined with prompts enriched of definition and guidelines, we name our approach SLIMER: Show Less, Instruct More - Entity Recognition [^1] https://huggingface.co/expertai/SLIMER.
Experiments were conducted on two standard NER benchmarks for zero-shot OOD, MIT (Liu et al., 2013) and CrossNER (Liu et al., 2021). Additionally, we assessed performance on never-seen-before NEs on BUSTER Zugarini et al. (2023), a document-level NER dataset with entity tags novel to all the evaluated models. We explored how increasing both the number of training samples and the number of unique Named Entity (NE) categories affect the generalization capabilities of LLMs, with or without the support of definitions and guidelines. Comparison of SLIMER with its baseline, i.e. the model devoid of definition and guidelines, reveals SLIMER’s deeper understanding, faster and more stable learning, and better zero-shot performance. Despite being trained on a fraction of the data, with little overlap between train and test named entity tags, SLIMER performs comparably against state-of-the-art instruction-tuned approaches, revealing stronger generalization capabilities when dealing with unseen named entities.
Model(s) MIT CrossNER TOT
@@
UniNER, GLiNER, GNER $10/12$ $/8$ 8 $/13$ 13 $10/11$ $/12$ 12 $/8$ 8 $/16$ 16 $52/55$
$83\%$ $100\%$ $100\%$ $91\%$ $100\%$ $100\%$ $100\%$ $95\%$
GoLLIE $1/12$ $3/8$ $5/13$ $5/11$ $5/12$ $6/8$ $7/16$ $12/55$
$8\%$ $38\%$ $38\%$ $45\%$ $42\%$ $75\%$ $44\%$ $22\%$
SLIMER $0/12$ $1/8$ $4/13$ $4/11$ $4/12$ $4/8$ $4/16$ $4/55$
$0\%$ $13\%$ $31\%$ $36\%$ $33\%$ $50\%$ $25\%$ $7\%$
Table 2: Overlap (%) between the NE types seen during training and evaluation by each model in MIT and CrossNER benchmarks. Since some classes are shared among different evaluation datasets, the final column accounts for the fraction relative to unique entity types. UniNER, GLINER, and GNER models exhibit near-complete overlap, meaning that in practice there are no unseen named entity classes in their test sets.
@@ 2 Related Work
Commonly employed machine-learning solutions frame NER into a “sequence labeling task”, where the goal is to assign a BIO label to each element in a given sequence Li et al. (2020). Fine-tuning BERT-family models (Devlin et al., 2018) for NER is a well established approach. While these models excel in supervised contexts, they have the severe limitation of being constrained to a predefined set of labels and inputs from limited domains, making it difficult for them to generalize across different contexts and on unseen named entities.
@@ 2.1 In-Context Learning
LLMs have demonstrated impressive few-shot and zero-shot capabilities on various challenging tasks Radford et al. (2019); Brown et al. (2020). However, endeavours to utilize LLMs for Information Extraction have been less promising Keraghel et al. (2024). Attempts to make use of LLMs (such as GPT) through clever prompt engineering have been conducted by Wang et al. (2023a) in their paper GPT-NER. In Ye et al. (2023), the authors compared several GPT models on various NLU tasks, including NER. The results highlighted a significant gap compared to supervised encoder-only approaches. Those results were further confirmed by Zhou et al. (2023) on other NER datasets.
@@ 2.2 Fine-Tuning for Zero-Shot NER
Parallel to In-Context Learning, other approaches explored Instruction-Tuning Wei et al. (2022); Chung et al. (2022); Wang et al. (2022b), focusing the training on a dedicated task, such as NER.
InstructUIE Wang et al. (2023b) is a encoder-decoder T5-11B model fine-tuned on supervised IE datasets (among which NER) phrased as text-to-text problems, the model is instructed to return a list of tuples (text span, entity type), choosing from a given list of categories. UniNER Zhou et al. (2023) consists in a decoder-only LLaMA model fine-tuned on a “conversation style template”. In inference, the model is prompted with the question “What describes NE in the text?” and a list of entities that belong to the requested NE is returned.
Several other works have followed, each based on a different instruction-tuning template. Table 1 provides an overview over the most significant state-of-the-art instruction-tuned models for NER, highlighting their architecture, the designed prompt template, and other key comparative features. GNER Ding et al. (2024) rethinks the importance of negative instances (i.e., “O” tags in BIO labeling) and replaces the established entity-centric schema with a BIO-like generation, replicating the same input text along with token-by-token BIO labels. Despite input length limitations and output parsing difficulties, their approach displays strong boundaries detection and reduced classification indecision. GLiNER Zaratiana et al. (2023) instead, relies on a much smaller, encoder-only and non-instruction-tuned architecture, which achieves remarkable performance in both supervised and zero-shot NER.
Definitions and guidelines for NER.
There have been some attempts to instruct models with additional information about the entity type(s). Zhou et al. (2023) experimented an UniNER variant enriched with brief entity definitions. However, such definitions were used in replacement of the entity name, rather than being provided as a supplementary aid. Generated contextually to a synthetic annotation process, they often varied for the same entity type, thus lacking of consistency, structure, and richness of content. As a result, UniNER-def model under-performed compared to its original version.
The effectiveness of adding NER tags descriptions was demonstrated in other works Sainz et al. (2024); Golde et al. (2024). In particular, GoLLIE’s authors Sainz et al. (2024), further explored this direction by adopting a code-based representation where the NE labels are encoded as Python classes, and guidelines are formatted as doc-strings. Comparison with a baseline excluding them demonstrated their effectiveness in tagging unseen NEs.
In this work, alike GoLLIE, we steer the model with annotation guidelines. However, we replace their code-based representation with a more natural language prompt that includes a definition and some guidelines for the category to be extracted. Our guidelines are longer, richer and more descriptive, yet simpler in syntax, making them a more intuitive and accessible solution for non-technical users who do not need to deal with code. Moreover, as also noted by GoLLIE’s authors, code-based syntax is more effective on code-oriented backbones. In contrast, these models are less familiar with long natural language sentences. This hinders the adoption of GoLLIE’s framework on a broad range of LLMs.
@@ 3 SLIMER
This section presents our approach, SLIMER. First, we provide motivations for reducing the number of training samples. Then, we describe the instruction-tuning prompt. Finally, we discuss how one can generate definitions and guidelines automatically by means of another LLM, such as ChatGPT.
@@ 3.3 Definition and Guidelines generation
To overcome this limitation, we exploited the OpenAI’s Chat-GPT APIs to automatically generate definition and guidelines. In particular, we designed the one-shot prompt template reported in Figure 2, which was used to query gpt-3.5-turbo-1106. An exemplary one-round user-system conversation is used to illustrate the desired output to the model. The three examples are randomly sampled for each NE from the dataset at hand. Thanks to such a prompt, all the generated definitions and guidelines exhibit a similar structure, with short and clear defining sentences, and with guidelines highlighting edge cases where to be cautious [^2] Some definition and guidelines examples in Appendix A..
@@ 4 Experiments
In the experiments, we investigate the impact of definition and guidelines on zero-shot NER. We compare SLIMER against state-of-the-art models on OOD inputs and unseen entity types.
@@ 4.1 Datasets
SLIMER models have all been fine-tuned on a subset of PileNER-type (Zhou et al., 2023), and evaluated on three different benchmarks: MIT (Liu et al., 2013), CrossNER (Liu et al., 2021) and BUSTER (Zugarini et al., 2023)
@@
PileNER-type (Zhou et al., 2023) is a synthetic dataset comprising a large set of approximately 50 thousands examples, encompassing over 13 thousands different named entity types. We kept only a subset of them, considering only those NEs with at least 100 instances. From the remaining 455 named entity classes, we manually revised and merged together classes of identical types spelled differently, e.g. organisation and organization, and discarded some “catch-all” gpt-hallucinated labels (e.g. unknown, other, miscellaneous, general, entity type), thus further reducing to 423 different labels. Finally, to limit the overlap between training and test entity types, we excluded nearly all the categories present in the test datasets, with the exception of the standard NER tags: person, location, organization, country. Overall, we kept 391 distinct NEs. Percentages of overlap between train and test are reported in Table 2. Further details about the dataset composition, grouped by topic, are illustrated in the appendix (see Figure 5).
@@
MIT (Liu et al., 2013) and CrossNER (Liu et al., 2021) datasets have become de-facto the standard benchmark for zero-shot NER. We use them to compare SLIMER against existing state-of-the-art models on out-of-domain (OOD) inputs.
@@
We extend the evaluation beyond the MIT and CrossNER benchmarks by including BUSTER (Zugarini et al., 2023). The dataset is a document-level NER benchmark in financial domain. Both domain and named entity tags differ from what observed by all the models during instruction-tuning, i.e. have 0% overlap. Its significant differences from standard NER datasets make BUSTER a perfect benchmark for evaluating zero-shot performance on never-seen-before NEs.
Model MIT CrossNER seen NEs
Movie Restaurant AI Literature Music Politics Science AVG
GPT-3.5-turbo 5.3 32.8 52.4 39.8 66.6 68.5 67.0 47.5 -
InstructUIE 63.0 21.0 49.0 47.2 53.2 48.2 49.3 47.3 20%
UniNER-type 42.4 31.7 53.5 59.4 65.0 60.8 61.1 53.4 95%
UniNER-def 27.1 27.9 44.5 49.2 55.8 57.5 52.9 45.0 $\geq$ 95%
UniNER-type+sup. 61.2 35.2 62.9 64.9 70.6 66.9 70.8 61.8 $\geq$ 95%
GoLLIE 63.0 43.4 59.1 62.7 67.8 57.2 55.5 58.4 22%
GLiNER-L 57.2 42.9 57.2 64.4 69.6 72.6 62.6 60.9 95%
GNER-T5 62.5 51.0 68.2 68.7 81.2 75.1 76.7 69.1 95%
GNER-LLaMA 68.6 47.5 63.1 68.2 75.7 69.4 69.9 66.1 95%
SLIMER w/o D&G $46.4\pm{1.8}$ $36.3\pm{2.1}$ $49.6\pm{3.2}$ $58.4\pm{1.7}$ $56.8\pm{2.1}$ $57.9\pm{2.1}$ $53.8\pm{1.7}$ $51.3\pm{2.0}$ 7%
SLIMER $50.9\pm 0.9$ $38.2\pm 0.3$ $50.1\pm 2.4$ $58.7\pm 0.2$ $60.0\pm 0.5$ $63.9\pm 1.0$ $56.3\pm 0.6$ $54.0\pm 0.5$ 7%
Table 3: Performance comparison between SLIMER and state-of-the-art models on the MIT and CrossNER benchmarks. Most models operate in an out-of-domain setting, due to their almost complete overlap between train and test tag sets. SLIMER instead tackles a more challenging scenario, with a significantly reduced train-test classes sharing. With the exception of UniNER-def, all the competitors’ results are taken from their respective papers as listed in Section 4.3.
Model Backbone Guidelines MIT CrossNER BUSTER
Movie Restaurant AI Literature Music Politics Science AVG
GNER LLaMA2-chat-7B $\times$ 37.1 34.3 47.1 50.4 60.4 52.7 55.3 48.2 14.4
GoLLIE LLaMA2-chat-7B $\checkmark$ 47.9 39.2 54.2 51.4 62.5 61.9 55.0 53.2 15.2
SLIMER LLaMA2-chat-7B $\checkmark$ 50.9 38.2 50.1 58.7 60.0 63.9 56.3 54.0 45.3
Table 4: Comparing SLIMER against state-of-the-art models when provided same LLM backbone and fine-tuning on the same sub-set of unique Named Entity types, such that the overlap between train-test sets is minimal.
@@ 4.2 Settings
Training details.
SLIMER is based on LLaMA-2 7B chat Touvron et al. (2023). Investigating how different families or model sizes affect results is outside the objectives of our work. In all the experiments the models were fine-tuned with LoRA (Hu et al., 2021) $r=8$, $\alpha=16$ for 10 epochs with early stopping, 32 batch size and learning rate initialized to $3.0\times 10^{-4}$ with cosine scheduler and a warm-up of 60 steps. Context length was set to 768, longer inputs were chunked. In order to reduce the overlap between training and test entities (see Table 2), we trained SLIMER on the PileNER-type subset described in Section 4.1, which comprises 391 distinct entity types. As already observed in literature, instruction-tuning helps aligning the model to the tasks, and most of the gains in performance are achievable with little amounts of instructions (Wang et al., 2022a; Zhou et al., 2024; Zugarini et al., 2024). Hence, we only picked 5 examples per class from the PileNER-type subset. In addition, for each example containing at least one annotated instance, we also included a negative example, i.e. one without any named entity. Overall, SLIMER was trained on 3910 input sequences, a small fraction of the data typically fed to other state-of-the-art models. To better understand the training set scales of different models, in Figure 3 we depicted them with circles proportional to their training sizes.